Find the first three terms of the arithmetic series in which asub1= 5 asubn= 33 and Ssubn= 152
Answer by stanbon(75887)

Find the first three terms of the arithmetic series in which asub1= 5 asubn= 33 and Ssubn= 152
-------------
a(1)=5
a(n)=a(1)+(n-1)d=33
5+(n-1)d=33
(n-1)d=28
------------
S(n)=(n/2)(a(1)+a(n))=152
(n/2)(5+33)=152
n*38=304
n=8
----------
Since n=8 and (n-1)d=28
7d=28
d=4
----------
So,
a(1)=5
a(2)=5+d=5+4=9
a(3)5+2d=5+8=13
-----------
Cheers,
Stan H.